# Principal stratification with $U$-statistics under principal ignorability

### v1

@@ 2.2 Principal generalized causal effects
Example 1 (Probability index under truncation by death).
@@
Example 2 (Win estimands under noncompliance).
@@ 3.1 Identification of the PGCE estimands
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome model models for specific contrast functions below.
Example 3 (Pairwise mean function for average treatment effect).
@@
Example 4 (Pairwise mean function for probability index).
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3. In this case, following Wolfe and Hogg, (1971), one can derive the pairwise outcome mean as
@@
Example 5 (Pairwise outcome mean for win estimands).
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao, (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 6.
Theorem 6 (Nonparametric identification).
@@
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 6, and the identification of the other two PGCEs follows the same strategy. In Theorem 6(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
@@
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by the Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 6(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
@@
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 6(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 6(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
Remark 2.
Theorem 6 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
@@ 3.2 Estimators based on nonparametric identification formulas
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart,, 1998). For example, Theorem 6(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
@@
Theorem 6(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
@@
and Theorem 6(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
@@ 4.1 The efficient influence function for the PGCE
The EIFs in Theorem 7 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ is then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
@@
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression, since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 7, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4.
@@ 4.2 Multiply robust estimators
The EIFs given in Theorem 7 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 7, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
@@
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 8.
Theorem 8 (Multiply robustness and efficiency under parametric working models).
Under Assumptions 13, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 8, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao, (2017) and Theorem 3 of Jiang et al., (2022). As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
Example 9 (Sample splitting for $U$-statistics).
@@
Theorem 10 (Efficiency under nonparametric working models).
Under Assumptions 13, and
@@
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 7.
Theorem 10 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 10 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Conditions (c) and (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al.,, 2018). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.

### v2

@@ 2.2 Principal generalized causal effects
Example 2.1 (Probability index under truncation by death).
@@
Example 2.2 (Win estimands under noncompliance).
@@ 3.1 Identification of the PGCE estimands
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome mean models for specific contrast functions below.
Example 3.1 (Pairwise mean function for average treatment effect).
@@
Example 3.2 (Pairwise mean function for probability index).
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.1. In this case, following Wolfe and Hogg, (1971), one can derive the pairwise outcome mean as
@@
Example 3.3 (Pairwise outcome mean for win estimands).
Consider the win estimands developed in Example 2.2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao, (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 3.4.
Theorem 3.4 (Nonparametric identification).
@@
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 3.4, and the identification of the other two PGCEs follows the same strategy. In Theorem 3.4(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
@@
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by the Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 3.4(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
@@
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 3.4(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 3.4(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
Remark 3.1.
Theorem 3.4 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
@@ 3.2 Estimators based on nonparametric identification formulas
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart,, 1998). For example, Theorem 3.4(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
@@
Theorem 3.4(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
@@
and Theorem 3.4(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
@@ 4.1 The efficient influence function for the PGCE
The EIFs in Theorem 4.1 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ is then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
@@
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression, since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 4.1, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4.1.
@@ 4.2 Multiply robust estimators
The EIFs given in Theorem 4.1 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 4.1, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
@@
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 4.2.
Theorem 4.2 (Multiply robustness and efficiency under parametric working models).
Under Assumptions 1-3, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 4.2, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao, (2017) and Theorem 3 of Jiang et al., (2022). As discussed in Remark 3.2, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
Example 4.3 (Sample splitting for $U$-statistics).
@@
Theorem 4.4 (Efficiency under nonparametric working models).
Under Assumptions 1-3, and
@@
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 4.1.
Theorem 4.4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 4.4 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Conditions (c) and (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al.,, 2018). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.